OpenStudy (anonymous):

Which of the following is a solution to the equation log 4 x + log 4 (x – 3) = 1 ?

OpenStudy (john_es):

Do you mean, \[\log_4x+\log_4(x-3)=1\]

OpenStudy (anonymous):

yes

OpenStudy (john_es):

Then use the properties of logarithms, \[\log_ax+\log_ay=\log_a(xy)\]\[\log_ax=1\Rightarrow x=a\] \[\log_4 x+\log_4(x-3)=1\Rightarrow \log_4\left(x(x-3)\right)=1\Rightarrow x(x-3)=4\] And solve for x. Do you understand it and know how to solve it?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

ohhh i get it now

OpenStudy (anonymous):

thanks

OpenStudy (john_es):

;), only to check it, the only (good) solution is 4.

OpenStudy (anonymous):

1 over the square root of 8 = 4(m + 3)

OpenStudy (anonymous):

can you help with this too

OpenStudy (john_es):

Do you mean, \[\frac{1}{\sqrt{8}}=4(m+3)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i taught it was 9/4 but its not

OpenStudy (john_es):

Then, it is a matter of solve for m, \[(m+3)=\frac{1}{4\sqrt{8}}\Rightarrow m=\frac{1}{4\sqrt{8}}-3\]

OpenStudy (john_es):

You can "simplify" a little of express it in various forms, \[m=\frac{1-24\sqrt{2}}{8\sqrt{2}}=\frac{\sqrt{2}-48}{16}\]

OpenStudy (john_es):

or express it, I mean.

OpenStudy (anonymous):

-2.911

OpenStudy (john_es):

Yes ;).